OpenStudy (anonymous):

48-13x3+[(50-7X5)+2]

OpenStudy (bahrom7893):

hmm this one would be: well should I just write this out as like the steps show.. in english?

OpenStudy (anonymous):

these confuse me with parenthsis

OpenStudy (anonymous):

p e m d a s

OpenStudy (bahrom7893):

forty eight minus thirteen times a number cubed plus the sum of fifty minus seven times a number to the fifth power and two

OpenStudy (bahrom7893):

u dont really need pemdas here..

OpenStudy (anonymous):

it sais simplify

OpenStudy (anonymous):

sorry

OpenStudy (bahrom7893):

ohhhhh haha sorry i thought this was also rewrite in english..

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

i was like what the hell

OpenStudy (bahrom7893):

48-13x3+[(50-7x5+2] 48-13x3+[52-7x5] 48+52-13x3-7x5 100-13x3-7x5

OpenStudy (bahrom7893):

that is as far as u can simplify..

OpenStudy (anonymous):

i thought it would be a number no? hmm?

OpenStudy (anonymous):

ok

OpenStudy (bahrom7893):

No it can't be a number.. the powers and coefficients of x are not the same.. if we had 7x5 instead of -13x3 the xs would cancel and we'd end up with 100 only..

OpenStudy (anonymous):

noooooooo thats times

OpenStudy (anonymous):

damn im fluttering up this morning

OpenStudy (anonymous):

your good though i need your brains

OpenStudy (bahrom7893):

ohhh i thought that was x.. please write times as *

OpenStudy (bahrom7893):

48-13x3+[(50-7X5)+2]= 48-39+(50-35+2) = 9+15+2= 26

OpenStudy (anonymous):

thank u!!!!

OpenStudy (bahrom7893):

you are very welcome... my head is hurtin.. i went to bed at 3am last nite... and woke up at 8.. ughh

OpenStudy (bahrom7893):

was watchin Inception.. that movie IS AMAZING!!!